Two years ago a woman was 4 times as old as her daughter. Three years from now the mother will be only three times as old as her daughter. How old are each of them now?
Answer by macston(5194)

.
M=mother's age; D=daughter's age
.
M-2=4(D-2)
M-2=4D-8
M=4D-6
.
M+3=3(D+3)
M+3=3D+9
M=3D+6
.
4D-6=3D+6
D-6=6
D=12
ANSWER 1: Daughter is 12 years old.
.
M=4D-6=4(12)-6=48-6=42
ANSWER 2: Mother is 42 years old.
.
CHECK:
M+3=3(D+3)
42+3=3(12+3)
45=3(15)
45=45
.